OpenStudy (jlg030597):

Ill give a medal and fan if you can solve it. There are four numbers given to you: 2, 3, 4 and 5. You are allowed to use only two mathematical symbols: + and = and come up with a mathematically true equation. Can you do it? PS- You can't repeat the symbols or numbers.

OpenStudy (jlg030597):

@AccessDenied @terryluce06 @ParthKohli @secret66 @Preetha

OpenStudy (anonymous):

Do you mean something like this? 2+5=3+4

OpenStudy (jlg030597):

Wrong the answer is 3^2=5+4